Diffusion Morning Report 8/20/03

ted berg tberg at snf.stanford.edu
Tue Aug 19 06:53:42 PDT 2003


Good Morning all,
    Everything looking good this morning.  Nitride investigation 
ongoing. Eleven hour run in progress.   What is going on with LTO it is 
still Yellow?


Diffusion 	Date: 	8/19/2003 	 
Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-Dirty Ox 	Up 	  	 
Tylan 8-SiGe 	Up 	  	9.9/30
Tylan 9-Poly 	Up 	  	17.6/30
Tylan 10-Nitride 	yellow 	Process testing 	9.9/30
Tylan 11-BPSG-LTO 	yellow 	 recipe issues 	26.1/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	9.3/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	  	 
Tylan 16 	N/A 	  	 
Burn Box 	Up 	  	 
ZYGO 	Up 	  	 
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	Auto changeover working well 	Currently 660- B(c7) 
on line



-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20030819/0c4f99ea/attachment.html>


More information about the diffusion mailing list